A franchise of automotive technicians has a contract arrangement for a rewards program with a car manufacturing company callePart (b)
Answer #1

Cost of services done by automotive technician = $23,000,000

Value of service for which car 1 car dollar issued = $5 of services

No of car dollar issued for the year =$23,000,000 / 5 = 46,00,000 car dollars

Value of car dollar issued by Automotive technician = $0.40

Stand alone selling price of car dollar by BMC = $0.45

Difference in value of car dollar i.e commission = 0.05

Commission = 46,00,000 * 0.05 = $2,30,000.00